OpenStudy (anonymous):

The equation, y = 10(1.0605)x, represents the winning prize money, in thousands, for first place in the Indianapolis 500 auto race. x represents the number of years after 1911, the first year. What is the value for 1911 (x = 0)? This is the prize money, in thousands. For instance, if you got 23, this represents $23,000.

OpenStudy (anonymous):

$10,605 $10,000

OpenStudy (misssunshinexxoxo):

y = 10(1.0605)^x, where x=0 y = 10(1.0605)^0 = 10(1) = 10 B. $10,000

OpenStudy (misssunshinexxoxo):

y = 10(1.0605)^x, where x = 50 y = 10(1.0605)^50 = 10(18.86) = 188.6 A. $189,000

OpenStudy (anonymous):

so B @misssunshinexxoxo

OpenStudy (misssunshinexxoxo):

You got it :) Way to go!

OpenStudy (anonymous):

ty